A 9 g lead bullet leaves the rifle at a temperature of 110°C and hits a steel plate. If the bullet melts on impact, what is the minimum speed it must have?
Answer #1

given
m = 9 g = 9*10^-3 kg

we know, specific heat of lead, C = 128 J/(kg C)

latent of melting of lead, L = 23*10^3 J/kg

we know, melting point of leat, T = 327.5 C

heat required to melt the lead bullet, Q = m*C*(327.5 - 110) + m*L

= 9*10^-3*128*(327.5 - 110) + 9*10^-3*23*10^3

= 457.6 J

so, reuired kinetic energy of the bullet, KE = 457.6 J

use, (1/2)*m*v^2 = KE

==> v = sqrt(2*KE/m)

= sqrt(2*457.6/(9*10^-3))

= 319 m/s <<<<<<<<<<<<<<<---------------------Answer